# Fourier Circuits in Neural Networks and Transformers: A Case Study of Modular Arithmetic with Multiple Inputs

### v3

@@ H.2 All Frequencies are Used for General $k$ Version
For this proof, for all $j_{1},\dots,j_{k+1}\in\mathbb{Z}$, to simplify the notation, let $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$, by Lemma F.8, so
@@
Setting aside the previously discussed points, it’s established in Lemma E.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
Appendix I Main Result
Section I.1 proves the main result for $k=3$. Section I.2 proves the general $k$ version of our main result.
I.1 Main result for $k=3$
Theorem I.1.
@@ I.1 Main result for $k=3$
By Lemma F.4, we get the single neuron class-weighted margin solution set $\Omega_{q}^{{}^{\prime}*}$ satisfying Condition 3.8 and $\gamma^{*}$.
By Lemma G.2 and Lemma E.1, we can construct network $\theta^{*}$ which uses neurons in $\Omega_{q}^{{}^{\prime}*}$ and satisfies Condition 3.8 and Definition 3.5 with respect to $q=\mathrm{unif}(\mathbb{Z}_{p})$. By Lemma E.2, we know it is the maximum-margin solution.
By Lemma H.1, when $j_{1}=j_{2}=j_{3}=-j_{4}\neq 0$, we must have $\widehat{f}(j_{1},j_{2},j_{3},j_{4})>0$. However, as discrete Fourier transform $\widehat{\phi}$ of each neuron is nonzero, for each frequency, we must have that there exists one neuron using it.
I.2 Main Result for General $k$ Version
Theorem I.2 (Formal version of Theorem 4.1).
@@ I.2 Main Result for General $k$ Version
Follow the same proof sketch as Theorem I.1 by Lemma F.8, Condition 3.8, Lemma G.3, Lemma E.1, Definition 3.5, Lemma E.2, Lemma H.2.
Appendix J More Empirical Details and Results
J.1 Implement Details
@@
J.2 One-hidden Layer Neural Network

### v4

@@ F.2 All Frequencies are Used for General $k$ Version
For this proof, for all $j_{1},\dots,j_{k+1}\in\mathbb{Z}$, to simplify the notation, let $\theta_{u}=\theta_{u}^{*}\cdot\frac{p}{2\pi}$, by Lemma D.8, so
@@
Setting aside the previously discussed points, it’s established in Lemma C.2 that the function $f$ maintains a consistent margin for various inputs as well as over different classes, i.e., $f$ can be broken down as
@@
Appendix G Proof of Main Result
Section G.1 proves the main result for $k=3$. Section G.2 proves the general $k$ version of our main result.
G.1 Main result for $k=3$
Theorem G.1.
@@ G.1 Main result for $k=3$
By Lemma D.4, we get the single neuron class-weighted margin solution set $\Omega_{q}^{{}^{\prime}*}$ satisfying Condition 3.8 and $\gamma^{*}$.
By Lemma E.2 and Lemma C.1, we can construct network $\theta^{*}$ which uses neurons in $\Omega_{q}^{{}^{\prime}*}$ and satisfies Condition 3.8 and Definition 3.5 with respect to $q=\mathrm{unif}(\mathbb{Z}_{p})$. By Lemma C.2, we know it is the maximum-margin solution.
By Lemma F.1, when $j_{1}=j_{2}=j_{3}=-j_{4}\neq 0$, we must have $\widehat{f}(j_{1},j_{2},j_{3},j_{4})>0$. However, as discrete Fourier transform $\widehat{\phi}$ of each neuron is nonzero, for each frequency, we must have that there exists one neuron using it.
G.2 Main Result for General $k$ Version
Theorem G.2 (Formal version of Theorem 4.1).
@@ G.2 Main Result for General $k$ Version
Follow the same proof sketch as Theorem G.1 by Lemma D.8, Condition 3.8, Lemma E.3, Lemma C.1, Definition 3.5, Lemma C.2, Lemma F.2.
Appendix H More Empirical Details and Results
H.1 Implement Details
@@
H.2 One-hidden Layer Neural Network